# Krylov Complexity for Open Quantum System: Dissipation and Decoherence

### v1

@@ 4.3 Decomposing the Master Equation into Individual Terms
Analzing the master equation it is clear that under particular operator choices and assumptions we can isolate the contributions from individual terms in the master equation and study how these terms contribute to Krylov complexity. Therefore in the following sections we will consider particular choices that allows us to investigate the individual contributions of decoherence and dissipation.
@@
This yields a master equation that accounts only for the coherent system dynamics and the environment-induced decoherence, as discussed in Ferialdi_2017 ; BRE02 ; Schlosshauer_2019 .
Figure 6: Left: Time evolution of Krylov complexity (K-complexity) for $\omega=1$ and $\gamma=1/100$. Right: Time evolution of Krylov complexity for the same parameters, with the decoherence function $\tilde{\Gamma}(t)$ overlaid in orange.
@@
For $\rho_{s}(t)$ in the form given by (41), it is evident that at least one of the diagonal elements of the reduced density matrix is time dependent, while at least one of the off-diagonal elements contains a term without time dependence. Consequently, there is no clear separation of the correlations carried by the off-diagonal terms, and the Krylov complexity cannot be reliably used to indicate the decay of coherences.
@@ 5 Discussion
In this work, we investigate Krylov complexity in open quantum systems, focusing on the damped harmonic oscillator and the Caldeira–Leggett model. Our results show that Krylov complexity effectively captures dissipative and decoherence effects, revealing distinct behaviors in these systems. In the Caldeira–Leggett model, the late-time Krylov complexity exhibits saturation, indicative of a fully mixed state driven by dissipation and decoherence, but it does not clearly reflect the onset of decoherence. In contrast, for the damped harmonic oscillator, where decoherence is suppressed, Krylov complexity highlights pronounced dissipative effects, characterized by a sharp initial increase followed by oscillatory decay. These findings underscore the potential of Krylov complexity as a diagnostic tool for probing open quantum systems and open avenues for further exploration in quantum information theory and quantum Brownian motion.
@@ Appendix A Krylov Complexity for the Harmonic Oscillator
$[H,[H,O]]$ $=\omega^{2}\alpha(a^{{\dagger}}+a)$ (50)
$,,$ (51)

### v2

@@ 4.3 Decomposing the Master Equation into Individual Terms
Analyzing the master equation, it is clear that under particular operator choices and assumptions, we can isolate the contributions from individual terms in the master equation and study how these terms contribute to Krylov complexity. Therefore, in the following sections, we will consider particular choices that allow us to investigate the individual contributions of decoherence and dissipation.
@@
This yields a master equation that accounts only for the coherent system dynamics and the environment-induced decoherence, as discussed in Ferialdi_2017; BRE02; Schlosshauer_2019. Using the same operator choice as in Sec. (4), we compute the two-point correlation function:
$\langle O(t)O\rangle_{\beta}$ $=e^{-2i\omega t}(C_{3}+3C_{2}+2C_{1})+e^{(-8m\gamma k_{B}T/\hbar^{2})t}(2C_{3}+3C_{2}+C_{1})$
$\quad+e^{(8m\gamma k_{B}T/\hbar^{2})t}(2C_{3}+C_{2})+e^{2i\omega t}(C_{3}-C_{2})$ (52)
Using the correlation function, we compute the K-complexity, which only consists of the coherent system dynamics and thermal decoherence term, as is displayed in Fig.(8) below:
Figure 8: Left: Time evolution of Krylov complexity (K-complexity) for $\omega=1$ and $\gamma=1/10$. Right: Time evolution of Krylov complexity for the same parameters, with the decoherence function $\tilde{\Gamma}(t)$ overlaid in orange.
@@
For $\rho_{s}(t)$ in the form given by (48), it is evident that at least one of the diagonal elements of the reduced density matrix is time dependent, while at least one of the off-diagonal elements contains a term without time dependence. Consequently, there is no clear separation of the correlations carried by the off-diagonal terms, and the Krylov complexity cannot be reliably used to indicate the decay of coherence.
@@ 5 Discussion
In this work, we investigate Krylov complexity in open quantum systems, focusing on the damped harmonic oscillator and the Caldeira–Leggett model. We emphasize that the choice of relatively simple models is deliberate, providing the analytical tractability (especially for evaluating correlation functions) essential to dissect the roles of dissipation and decoherence via Krylov techniques. Although Krylov complexity and associated Lanczos coefficients have traditionally been leveraged to explore quantum chaos, this study re-purposes them to pinpoint these effects in open quantum systems. As far as we are aware, ours is an early effort to deploy Krylov methods for this purpose, shedding new light on the intricate interplay between dissipation and decoherence. Our results show that Krylov complexity effectively captures dissipative and decoherence effects, revealing distinct behaviors in these systems. In the Caldeira–Leggett model, the late-time Krylov complexity exhibits saturation, indicative of a fully mixed state driven by dissipation and decoherence, but it does not clearly reflect the onset of decoherence. In contrast, for the damped harmonic oscillator, where decoherence is suppressed, Krylov complexity highlights pronounced dissipative effects, characterized by a sharp initial increase followed by oscillatory decay. These findings underscore the potential of Krylov complexity as a diagnostic tool for probing open quantum systems and open avenues for further exploration in quantum information theory and quantum Brownian motion.
@@ Appendix A Krylov Complexity for the Harmonic Oscillator
$[H,[H,O]]$ $=\omega^{2}\alpha(a^{{\dagger}}+a)\,$, (57)
@@
Appendix B Correlation functions and Time evolved operators
Here we show the general procedure followed to compute the time-evolved operators for the Caldeira-Leggett master equation in Sec. (4). We only show the procedure followed in the first case, since for the rest of the operators, only a slight change is needed. We begin with the full form of the Lindblad master equation as given in (34), with the initial operator being $O(0)=p^{2}$, this is because it is the simplest form of operator that allows us to obtain contributions from both the dissipative and decoherence terms. The Lindblad jump operator $L$ is given in (35). The general procedure we will follow here is similar to Sec. (3). The evolution of such operators when written in terms of creation and annihilation operators can, in general, be written in terms of some time-dependent functions $f_{i}(t)\in\mathbb{C}$. As such, initially, we will leave the function, as well as the time dependence, implicit, focusing only on the operator algebra, and reintroduce the time dependence later. We begin by considering the individual contributions from each term of (34). From the first term, for the coherent dynamics, we have:
$i[H_{s},p^{2}]$ $=i\left[\left(\frac{p^{2}}{2}+\frac{\omega^{2}x^{2}}{2}\right),p^{2}\right]\,$,
$=i\omega^{2}\left[\frac{x^{2}}{2},p^{2}\right]\,$,
$=i\omega^{2}\left[\frac{x^{2}}{2},p\right]p+i\omega^{2}p\left[\frac{x^{2}}{2},p\right]\,$,
$=i\omega^{2}(ixp+ipx)\,$,
$=-\omega^{2}xp-\omega^{2}px\,.$ (64)
The second term of (34), which behaves like a Hamiltonian contribution and comes from the dissipator, yields the following:
$i\frac{\gamma}{2}[xp+px,p^{2}]$ $=i\frac{\gamma}{2}[xp,p^{2}]+i\frac{\gamma}{2}[px,p^{2}]\,$,
$=i\frac{\gamma}{2}[2ip^{2}]+i\frac{\gamma}{2}[2ip^{2}]\,$,
$=-2\gamma p^{2}\,.$ (65)
Now we tackle the contributions which come from the dissipator in the Lindblad master equation, the third term of the master equation (34) gives us:
$L^{\dagger}p^{2}L$ $=(Ax-iBp)(p^{2})(Ax+iBp)\,$,
$=A^{2}xp^{2}x+iABxp^{3}+B^{2}p^{4}-iABp^{3}x\,.$ (66)
From the fourth term (of (34)), we get:
$-\frac{1}{2}L^{\dagger}Lp^{2}$ $=(Ax-iBp)(Ax+iBp)(p^{2})\,$,
$=-\frac{1}{2}(A^{2}x^{2}p^{2}+iABxp^{3}-iABpxp^{2}+B^{2}p^{4})$ (67)
and finally, from the fifth term of (34) we get:
$-\frac{1}{2}p^{2}L^{\dagger}L$ $=(p^{2})(Ax-iBp)(Ax+iBp)\,$,
$=-\frac{1}{2}(A^{2}p^{2}x^{2}+iABp^{2}xp-iABp^{3}x+B^{2}p^{4})\,.$ (68)
Combining all of them, we get,
$S$ $=\gamma(A^{2}xp^{2}x-\frac{1}{2}A^{2}x^{2}p^{2}-\frac{1}{2}A^{2}p^{2}x^{2})+\gamma(\frac{i}{2}ABxp^{3}-\frac{i}{2}ABp^{3}x)+\gamma(\frac{i}{2}ABpxp^{2}-\frac{i}{2}ABp^{2}xp)\,$,
$=\gamma(A^{2}xp^{2}x-\frac{1}{2}A^{2}x^{2}p^{2}-\frac{1}{2}A^{2}p^{2}x^{2})+\gamma\frac{iAB}{2}[x,p^{3}]+\gamma\frac{iAB}{2}[pxp,p]\,$,
$=\gamma(A^{2}xp^{2}x-\frac{1}{2}A^{2}x^{2}p^{2}+[x^{2},p^{2}]-\frac{1}{2}A^{2}x^{2}p^{2})+\gamma\frac{iAB}{2}[x,p^{3}]+\gamma\frac{iAB}{2}[pxp,p]\,$,
$=\gamma(A^{2}xp^{2}x-A^{2}x^{2}p^{2}+[x^{2},p^{2}])+\gamma\frac{iAB}{2}[x,p^{3}]+\gamma\frac{iAB}{2}[pxp,p]\,$,
$=\gamma(A^{2}x[p^{2},x]+[x^{2},p^{2}])+\gamma\frac{iAB}{2}[x,p^{3}]+\gamma\frac{iAB}{2}[pxp,p]\,$,
$=\gamma(A^{2}-2ABp^{2})\,$, (69)
This leaves us with the following:
$\frac{d}{dt}O$ $\propto-\omega^{2}xp-\omega^{2}px-2\gamma p^{2}+\gamma(A^{2}-2ABp^{2})\,$,
$\propto-\omega^{2}xp-\omega^{2}px+(-2\gamma-\gamma 2AB)p^{2}+\gamma A^{2}(aa^{\dagger}-a^{\dagger}a)\,.$ (70)
This is the equation (38) that we have used in the main text. Now to solve for the $O(t),$ we first recall that $a(t)=f^{*}(t)a$ and similarly $a^{{\dagger}}(t)=f(t)a^{{\dagger}}$ as discussed in Sec. (3). We then rewrite every term of (70) in terms of $a(t)$ and $a^{{\dagger}}(t)$ in the following way,
$(-\omega^{2})xp$ $=\underbrace{(-\frac{i}{2}\omega^{2})}_{M}(a^{\dagger}(t)a(t)-a^{\dagger}(t)a^{\dagger}(t)+a(t)a(t)-a(t)a(t)^{\dagger})\,$,
$(-\omega^{2})p\,x$ $=\underbrace{(-\frac{i}{2}\omega^{2})}_{N}(a(t)a^{\dagger}(t)-a^{\dagger}(t)a^{\dagger}(t)+a(t)a(t)-a^{\dagger}(t)a(t))\,$,
$(-2\gamma-2\gamma AB)p^{2}$ $=\underbrace{\frac{\omega}{2}(2\gamma+2\gamma AB)}_{O}(a(t)a(t)-a(t)a^{\dagger}(t)-a^{\dagger}(t)a(t)+a^{\dagger}(t)a^{\dagger}(t))\,.$ (71)
We now have the following form of the differential equation.
$\frac{d}{dt}O$ $=-(-M+N+O)f_{1}(t)a^{\dagger}a+(-M-N-O)f_{2}(t)a^{\dagger}a^{\dagger}$
$+(M+N+O)f_{3}(t)aa-(M-N+O)f_{4}(t)aa^{\dagger}\,.$ (72)
The functions $f_{1},f_{2},f_{3},f_{4}$ can be read off from (28). Finally, we get an explicit form of the time-evolved operator as:
$O(t)$ $=-\frac{\omega}{2}(\exp[(2i\omega-2\gamma-2\gamma AB)t]aa$
$-\exp[(\gamma A^{2}-2\gamma-2\gamma AB)t]aa^{\dagger}$
$-\exp[(-\gamma A^{2}-2\gamma-2\gamma AB)t]a^{\dagger}a$
$+\exp[(-2i\omega-2\gamma-2\gamma AB)t]a^{\dagger}a^{\dagger})\,.$ (73)
Equipped with this, we can therefore compute the correlator:
$\langle O(t)O\rangle_{\beta}$ $=\frac{\mathrm{Tr}\{\rho_{\beta}O(t)O\}}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\,$,
$=\frac{1}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\sum_{n}e^{-\beta(n+1/2)\omega}\langle n|\frac{\omega^{2}}{4}(e^{(2i\omega-2\gamma-2\gamma AB)t}aa-e^{(\gamma A^{2}-2\gamma-2\gamma AB)t}aa^{\dagger}$
$-e^{(-\gamma A^{2}-2\gamma-2\gamma AB)t}a^{\dagger}a+e^{(-2i\omega-2\gamma-2\gamma AB)t}a^{\dagger}a^{\dagger})(a^{{\dagger}}a^{{\dagger}}-a^{{\dagger}}a-aa^{{\dagger}}+aa)|n\rangle\,$,
$=\frac{1}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\frac{\omega^{2}}{4}\sum_{n}e^{-\beta(n+1/2)\omega}(e^{(2i\omega-2\gamma-2\gamma AB)t}(n^{2}+3n+2)$
$+e^{(\gamma A^{2}-2\gamma-2\gamma AB)t}(2n^{2}+3n+1)+e^{(-\gamma A^{2}-2\gamma-2\gamma AB)t}(2n^{2}+n)$
$+e^{(-2i\omega-2\gamma-2\gamma AB)t}(n^{2}-n))\,.$ (74)
Here, the choice of normalization is the common choice for normalization to ensure that at C(0)=1.C(0)=1\,. Now we use the following to perform the sum over $n$,
$\begin{split}&\sum_{n}\langle n|a^{{\dagger}}a^{{\dagger}}aa|n\rangle=n^{2}-n\,,\quad\sum_{n}\langle n|aaa^{{\dagger}}a^{{\dagger}}|n\rangle=n^{2}+3n+2\,,\\ &\sum_{n}\langle n|aa^{{\dagger}}aa^{{\dagger}}|n\rangle=n^{2}+2n+1\,,\quad\sum_{n}\langle n|aa^{{\dagger}}a^{{\dagger}}a|n\rangle=n^{2}+n\,,\\ &\sum_{n}\langle n|a^{{\dagger}}aa^{{\dagger}}a|n\rangle=n^{2}\,,\quad\sum_{n}\langle n|a^{{\dagger}}aaa^{{\dagger}}|n\rangle=n^{2}+n\,.\end{split}$
Finally, one can use the expressions for $C_{1},C_{2},C_{3}$ as mentioned in (42) of Sec . 4 to bring it into a more compact form. The procedure outlined above was used in computing the operators $O(t)$ and correlators for the entirety of Sec. (4).